Calculate the higher heating value in kJ/mand kJ/kg at 10 °C and 3 atm for an LPG gas mixture 60% propane and 40% butane
Answer #1

Given conditon :

100C and 3 atm

60% propane and 40% butane

The higher heating values (HHVs) of the hydrocarbon compounds can be calculated based on the carbon (C) and hydrogen (H) contents of the chemical structures. HHVs (MJ / kg) as a function of the carbon (C) and hydrogen (H) fractions of N-saturated hydrocarbons can be calculated by the following equation:

Higher heating value (HHV) = 0.303 (C) + 1.423 (H) MJ/kg

by using the above formula, HHV is calculated

Propane (C3H8) = 50.22 MJ/Kg

Butane (c4H10) = 49.2 MJ/Kg

Given 60% propane and 40% butane,

HHV = (0.60*50.22) + (0.40*49.2) = 49.812 MJ/Kg = 49812 KJ/Kg

density of propane = 493 Kg/m3 at 100C

density of butane = 573 Kg/m3 at 100C

for propane HHV = 50.22* 493 = 24758.46 MJ/m3

for propane HHV = 49.2* 573 = 28191.6 MJ/m3

Given 60% propane and 40% butane,

HHV = (0.60*24758.46) + (0.40*28191.6) = 26131.716 MJ/m3 = 26131716 KJ/m3